A university student painter is considering the purchase of a new air compressor and paint gun to replace an old paint sprayer. (Both items belong to Class 9 and have a 25 percent CCA rate.) These two new items cost $12,000 and have a useful life of four years, at which time they can be sold for $1,600. The old paint sprayer can be sold now for $500 and could be scrapped for $250 in four years. The entrepreneurial student believes that operating revenues will increase annually by $8,000. The tax rate is 22 percent and the required rate of return is 15 percent.

Find the PV for the following: initial cost, after-tax operating cash flows, CCA tac shield, salvage, loss of tax shield due to salvage.

Find NPV
Answer #1

Present value of initial cost = (12000-500)-((1600-250)/(1.15^4)) = $10728.13

Present value of operating income = 8000*(1-22%)*PVIFA(15%,) = $17815.06

Present value of (CCA TS OLD) = (((500*25%*15%)/(25%+15%))+((1.5*15%)/(1.15)))- (((250*25%*15%)/(25%+15%))+(1/1.15^4)) = $23.06

Present value of (CCA TS NEW) = (((12000*25%*15%)/(25%+15%))+((1.5*15%)/(1.15)))- (((1600*25%*15%)/(25%+15%))+(1/1.15^4)) = $974.62

NPV = present value of cash inflows – present value cash out flows

= (17815.06+(974.62-23.06))- 10728.13 = $8038.49

As NPV is positive, the new air compressor and paint gun should be purchased.